OpenStudy (anonymous):

65 concert ticket;7.5 sales tax?

OpenStudy (anonymous):

Change 7.5% into a decimal.

OpenStudy (anonymous):

Do you know how? @foofoo787

OpenStudy (anonymous):

oh yes is it 0.075?

OpenStudy (anonymous):

Yes, now multiply 65 by the decimal of 7.5%

OpenStudy (anonymous):

4.875

OpenStudy (anonymous):

Yes, now add 65 and that number.(Round the number to the nearest tenth)

OpenStudy (anonymous):

69.875

OpenStudy (anonymous):

69.88?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

:'D THANK YOUUUUUUU i really wanna hug you

OpenStudy (mathstudent55):

Correct. You're getting the hang of it.

OpenStudy (anonymous):

no problem :)

OpenStudy (mathstudent55):

Remember these things: 1) To find a percentage of a number, change the percent to a decimal and multiply by the number. 2) If it's a tax, add that percentage to the number. 3) If it's a discount, subtract the percentage from the number.

OpenStudy (anonymous):

oh. I think that was my problem. I didn't know if i should add or subtract

OpenStudy (anonymous):

Here is a sample problem. Try to do it by yourself first. 26% off of $112.25

OpenStudy (anonymous):

ok:)

OpenStudy (mathstudent55):

Example 1. $50 book with 6% tax. 6% * $50 = 0.06 * 50 = $3 (this is the tax) Final price is $50 + $3 = $53 (price of book plus tax). Example 2. $20 ticket with 8% discount. 8% * $20 = 0.08 * 20 = $1.60 (this is amount of discount) $20 - $1.60 = $18.40 (this is the final price after taking off the discount)